OpenStudy (anonymous):

A paddle boat can move at a speed of 10 km/h in still water. The boat is paddled 6km downstream in a river in the same time it takes to go 3km upstream. What is the speed of the river?

OpenStudy (kira_yamato):

Is the speed of the boat constant at 10km/h?

OpenStudy (anonymous):

That is all the info they give me.

OpenStudy (anonymous):

What do you mean @Chlorophyll?

OpenStudy (anonymous):

Downstream: ( 10 + r ) t = 6 km Upstream: ( 10 - r ) t = 3 km

OpenStudy (anonymous):

Hint: equate them ( 10 + r ) / 6 = ( 10 -r ) /3

OpenStudy (anonymous):

\[\frac{ 6 }{ 10+x }=\frac{ 3 }{ 10-x }\]

OpenStudy (anonymous):

Right I got that but for some reason I am still not coming up with an answer that looks correct. I got 10/3

OpenStudy (anonymous):

Well that's what I got was x=10/3

OpenStudy (anonymous):

Well the computer accepted that... I don't know what I was doing wrong... I kept just getting 10.

OpenStudy (anonymous):

Thank you!